# Conceptualization, Writing—original draft

### v2

@@ 1 Introduction
Edge detection captures object boundaries and salient edges in images, which can keep the important semantic information in the image and discard the irrelevant noise such as color and texture. As a result, edge detection is one of the most important basic tasks of computer vision and is widely used in many downstream tasks, such as image segmentation [20], depth map prediction [44] and salient detection [42].
Early researches on edge detection [3, 26, 29, 2, 23] mainly rely on local features, but the local features of edges and texture noise are highly similar and hard to distinguish. Therefore, the performance of early edge detectors is hardly satisfactory. With the application of deep learning in edge detection [36, 18, 11, 13, 30], the performance of detectors has been improved unprecedented, far surpassing the local feature based methods and even surpassing human beings. Deep learning based edge detectors have become the mainstream methods for current edge detection.
As the pioneer of contemporary edge detection, HED first introduces the deep supervision mechanism to edge detection and learns multi-scale predictions holistically, which dominates the development of edge detection at that time [18, 11, 13]. To keep the lightweight of the model, these methods use extremely simple decoders to fuse and compress the multi-scale information into single-channel edges. With the development of deep learning, the encoder and decoder are gradually upgraded iteratively. Puet al. [24] replace the encoder with Transformer. Denget al. [6] introduce encoder-decoder with symmetric structure into edge detection. But this encoder-decoder structure has remained the same. The encoder extracts multi-scale features including location information and semantic information, and the decoder fuses these features to obtain the final edge map.
In recent years, the development of this structure seems to reach a bottleneck and researchers turn to focus on the optimization of data processing and training strategies [43, 40, 17]. These methods further improve the performance of edge detectors by using complex data processing and training strategies such as data distillation [17], multi-stage training [24], multi-step inference [40], auxiliary model [8], etc.
@@ 2.1 The Encoder of Edge Detection
The pioneer of contemporary edge detection is HED [36], which extracts features using VGG [28], the state-of-the-art feature extraction network at the time, and fuses multi-scale features with several convolutions and upsampling to get the final edge. Although rough, HED fuses local and global information for the first time, which makes it perform much better than other detectors based on local information and achieves the highest accuracy. Inspired by HED, a series of excellent edge detectors [18, 11, 13] have emerged. For fair comparison, there is a tacit agreement to use VGG to extract features and make the decoder as lightweight as possible.
With the development of deep learning, stubborn selection of VGG to extract features is not only not conducive to further improvement of model accuracy, but also difficult to flexibly adjust the size of the model. Therefore, researchers begin to explore new encoders. On the one hand, they explore more powerful encoders to obtain higher accuracy edges. Both RCF [18] and BDCN [11] try to directly use ResNet [12], a more advanced encoder to improve the performance of their models. EDTER [24] introduces Transformer into edge detection for the first time, which brings huge computation, but also improves the accuracy by leaps and bounds. The important innovation of DiffusionEdge [40] is to introduce Diffusion to edge detection, but its encoder is also updated with the latest SWin-Transformer. On the other hand, more efficient encoders are designed to obtain lightweight models, FINED [35] only uses the first few convolutions of VGG to avoid a lot of computation, and LDC [30] utilizes channel separable convolutions and a smaller number of channels to design a more lightweight encoder. PiDiNet [32] combines channel separable convolutions with handcrafted features to obtain high-quality edges with a lighter weight model.
Edge detection requires location information and semantic information. In previous works [18, 24, 43], both are extracted by the same encoder. This may seem to simplify the design of the model, but shallow features need to maintain high resolution to ensure the accuracy of location information, while keep a sufficient number of channels to provide rich information cues for deep features. This leads to an inevitable increase in model parameters and computational cost. This problem is exacerbated after Transformer is introduced to edge detection [24, 40], which results in tens of times more parameters compared to previous CNN-based methods [18, 11].
@@ 2.2 The Decoder of Edge Detection
To the best of our knowledge, the two most commonly used decoders in edge detection are the HED-style decoder [36, 11, 38] and the UNet-style decoder [6, 4]. The HED-style decoder fuses and compresses the multi-scale features into a single-channel edge map using only simple upsampling and $1\times 1$ convolution. The HED-style decoder can obtain the location information in the shallow features and the semantic information in the deep features at the same time. However, shallow features will contain a lot of noise due to the lack of semantic information, and semantic features will introduce location errors due to excessive downsampling. The existence of these two kinds of errors seriously affects the quality of the generated edge map. The UNet-style decoder gradually fuses high-level semantic information with shallow features, and compresses the fused shallowest features into a single-channel edge map. Therefore, the adverse effects of location errors in the semantic information on the edge map can be alleviated. However, the close contact between shallow features and the final result inevitably mixes noise into the generated edge map.
@@ 2.3 Training Strategies for edge detection
UAED [43] studies the subjectivity and ambiguity of different annotations by estimating the uncertainty of the data. This not only requires the dataset to have multiple sets of labels, but also requires additional processing of these labels. KDED [17] and PEdge [8] eliminate the adverse effects caused by subjectivity and ambiguity of annotations by distillation and self-distillation, which make the training process more complex. PiDiNet [32] and PiDiNeXt [16] achieve efficient inference through reparameterization, and EDTER [24] achieves global context modeling and local refinement through multi-stage training, respectively. DiffusionEdge [40] introduces Diffusion to the edge detection task and transforms edge detection into a multi-step denoising process.
@@ 3 A NEW BASELINE FOR EDGE DETECTION
Figure 1: (a) Overall framework of NBED. SeparConvs and Attentions contain multiple SeparConv and Attention modules, respectively. n×$\times$ down means n $\times$ downsampling. $E_{i}$: Feature size at stage i of the encoder, $D_{i}$: Feature size at stage i of the decoder. (b) Meta block. SeparConv and Attention in (a) are composed of Meta block and different Operators.
@@ 3.1 Location Feature Encoder
In previous works [18, 11, 38], multi-scale features are usually extracted by the same encoder. As a result, the shallow module is not only used to generate location features, but also used to provide cues for higher-level modules. For the shallow feature $L_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$, to ensure the accuracy of the location information, the resolution $H_{i}\times W_{i}$ must be large enough. At the same time, the number of channels $C_{i}$ cannot be reduced to keep rich cues for semantic information. So $L_{i}$ is usually a feature with heavy parameters. In addition, the problem of gradient instability forces researchers to employ additional processing such as regularization [15] and residual connection [12], even in shallow modules. It makes the computation of shallow features more complex.
We propose an independent location feature encoder, a shallow network. The location feature encoder contains two vanilla Convolution blocks, and each vanilla convolution block consists of a $3\times 3$ convolution module and a ReLU module. To ensure the accuracy of the location information, the feature resolution in the first vanilla Convolution block is consistent with the input image. A down-sampling module is inserted between the two convolution blocks to generate two-scale location features, which can reduce the scale gap between location features and semantic features. Consistent with VGG16 [28], the number of channels doubles when the feature resolution is halved. Since providing cues for semantic information is no longer considered, the features in Location Feature Encoder is compressed to 16 channels, which is much smaller than existing models such as VGG16 [28] (64 channels) and ResNet50 [12] (64 channels). Compared to VGG16 and ResNet50, our Location Feature Encoder not only has fewer modules, but also has fewer channels per module. Therefore, even though the features keep high resolution, the computational cost remains low.
@@ 3.2 Semantic Feature Encoder
The cascaded residual block is actually an important component of the Transformer and together with the Self-Attention module contributes to the success of the Transformer [21]. We intended to use Transformer to capture the semantic features needed for edge detection, but the massive computational cost of Self-Attention (SA) prevents us, even if the image is downsampled before being fed into the semantic feature encoder. Inspired by CAFormer [41], we adopt a CNN-Transformer hybrid model that uses CNN in the first two stages with large resolution and SA only on the low-resolution features in the last stage. To further reduce the computational cost, the CNN in the first two stages is replaced by channel separable convolution, which is consistent with CAFormer [41].
@@ 3.4 Network Training
We adopt the annotator-robust loss proposed in RCF [18] for all generated edge maps. It can be formulated as
@@
where p denotes the final edge prediction and y is the ground truth. $\alpha=\lvert Y_{+}\rvert/\lvert Y\rvert$ and $\beta=\lambda\cdot\lvert Y_{-}\rvert/\lvert Y\rvert$. $\lvert Y_{+}\rvert$ and $\lvert Y_{-}\rvert$ are used to represents the number of edge and non-edge, separately. $\lambda$ controls the weight of positive over negative samples. $\lvert Y\rvert=\lvert Y_{+}\rvert+\lvert Y_{-}\rvert$. Due to the inconsistency of annotations among different annotators, a threshold $\gamma$ is introduced for loss computation [31]. For pixel i, it will be regarded as edge if the ground truth $y_{i}$ is more than $\gamma$, and we define the ground truth $y_{i}=1$. Pixel i will be regarded as no-edge if $y_{i}=0$. $\lambda$ controls the weight of edge over non-edge.
@@ 4.1 Datasets
We evaluate the proposed NBED on three widely used datasets BSDS500, NYUDv2, and BIPED. BSDS500 [1] is by far the most widely used dataset in edge detection, which includes 200 training images, 100 validation images, and 200 test images. Four to nine annotated maps are associated with each image. The final edge ground truth is computed by taking their average. NBED are trained on training and validation images, and tested on the test images. NYUDv2 [27] is an indoor scene semantic segmentation dataset, the edge ground truth is generated from segmentation maps. it contains 1449 groups of carefully annotated RGB and depth images, and each group image has one annotation result. We use 795 images to train the model and evaluate the model on the rest images. BIPED [22, 31] contains 250 carefully annotated high-resolution Barcelona Street View images. There are 200 images for training and validation, and 50 images for testing. All images are carefully annotated at single-pixel width by experts in the computer vision field.
The data augmentation strategy during training is shown in Table 2. The data augmentation strategy of BSDS is consistent with LPCB [6]. The data augmentation strategy of NYUDv2 follows RCF [18], with the only difference that the augmentation is randomly cropped to $400\times 400$ for batch calculation. The data augmentation strategy of BIPED is consistent with DexiNed [22].
@@
BSDS F (4×4$\times$4 ×), R (25×25$\times$25 ×) $481\times 321$
NYUDv2 F (2×2$\times$2 ×), S (3×3$\times$3 ×), R (4×4$\times$4 ×) $400\times 400$
BIPED F (2×2$\times$2 ×), S (3×3$\times$3 ×), R (16×16$\times$16 ×), G (3×3$\times$3 ×) $400\times 400$
@@ 4.2 Implementation details
Due to local correlation, the edges obtained by deep learning-based methods are edge probability maps, even after NMS processing. So we need to choose a threshold to binarize the edges. There are two options for the threshold used to binarize the edges. One is to find an optimal threshold value for each image (OIS), and another is to use the same optimal threshold for the whole dataset (ODS). The maximum tolerance allowed for correct matches between edge predictions and ground truth annotations is set to 0.0075 for BSDS500 and BIPED, and 0.011 for NYUDv2. For more experimental details, please refer to previous works [13, 22].
@@ 4.3 Comparison with state of the arts
Methods Pub.’Year SS MS Param. (M) GFLOPs
@@
Canny [2] PAMI’86 0.611 0.676 - - - -
gPb-UCM [1] PAMI’10 0.729 0.755 - - - -
SCG [25] NeurIPS’12 0.739 0.758 - - - -
SE [7] PAMI’14 0.743 0.764 - - - -
OEF [10] CVPR’15 0.746 0.770 - - - -
DeepEdge [9] CVPR’15 0.753 0.772 - - - -
DeepContour [34] CVPR’15 0.757 0.776 - - - -
Deep Boundary [14] ICLR’15 0.789 0.811 0.803 0.820 - -
CEDN [39] CVPR’16 0.788 0.804 - - - -
RDS [19] CVPR’16 0.792 0.810 - - - -
AMH-Net [37] NeurIPS’17 0.798 0.829 - - - -
HED [36] PAMI’17 0.788 0.808 - - 14.6 57.5
CED [33] CVPR’17 0.803 0.820 - - 21.8 -
LPCB [6] ECCV’18 0.800 0.816 - - - -
RCF [18] PAMI’19 0.798 0.815 - - 14.8 75.3
DSDC [4] ACMMM’20 0.802 0.817 - - - -
LDC [5] ACMMM’21 0.799 0.816 - - - -
PiDiNet [32] ICCV’21 0.789 0.803 - - 0.71 11.6
BDCN [11] PAMI’22 0.806 0.826 - - 16.3 103.4
FCL [38] NN’22 0.807 0.822 0.816 0.833 16.5 134.4
EDTER [24] CVPR’22 0.824 0.841 0.840 0.858 468.8 802.3
UAED [43] CVPR’23 0.829 0.847 0.837 0.855 72.5 72.8
DiffusionEdge [40] AAAI’24 0.834 0.848 - - 225 494
@@
We compare our model with traditional detectors (Canny [2], SCG [25], OEF [10], etc), CNN-based detectors (RCF [18], BDCN [11], UAED [43], etc) and the transformer-based detectors (EDTER [24] and DiffusionEdge [40]) on BSDS500. Quantitative results are shown in Table 3. By training on the trainval set of BSDS500, our NBED achieves the F-measure ODS of 0.834 with single-scale testing and obtains 0.840 with multi-scale testing, which is superior to all the state-of-the-art edge detectors.
It’s worth mentioning that NBED is a vanilla encoder-decoder with neither a complex training strategy nor a huge model size, which stands out among recent high performance edge detectors. Among the detectors with similar performance to our NBED, EDTER [24] not only has a model size of 11 times larger than NBED, but also requires multi-stage training. Nonetheless, the single-scale ODS of EDTER is still 1.4% lower than that of NBED. The number of parameters of UAED [43] is about twice that of NBED, and it requires multi-label data. The single-scale ODS of UAED in this case is 0.9% lower than that of NBED. DiffusionEdge [40] appears to perform competitively, with a single-scale ODS 0.4% lower than NBED’s, and DiffusionEdge not only relies on large-scale models, but also requires multi-step sampling during inference, which is very inefficient. In addition to these methods with complex training strategies, the top performing vanilla encoder-decoder is FCL [38] with a single-scale ODS of 0.807, which is more than 3% lower than NBED and is no longer a contender. Some qualitative results are shown in Fig. 2. We observe that the proposed NBED shows a clear advantage in prediction quality, both crisp and accurate.

### v3

@@ Abstract
Note: This work has been accepted for publication in Signal Processing: Image Communication. A definitive version is available at: https://doi.org/10.1016/j.image.2026.117485
@@ 1 Introduction
Edge detection captures object boundaries and salient edges in images, which can keep the important semantic information in the image and discard the irrelevant noise such as color and texture. As a result, edge detection is one of the most important basic tasks of computer vision and is widely used in many downstream tasks, such as image segmentation [liu2021dance], depth map prediction [zhu2020edge] and salient detection [zhao2019egnet].
Early researches on edge detection [davis1975survey, sharifi2002classified, sobel19683x3, canny1986computational, prewitt1970object] mainly rely on local features, but the local features of edges and texture noise are highly similar and hard to distinguish. Therefore, the performance of early edge detectors is hardly satisfactory. With the application of deep learning in edge detection [xie2017holistically, liu2019richer, he2022bdcn, huan2021unmixing, soria2022ldc], the performance of detectors has been improved unprecedented, far surpassing the local feature based methods and even surpassing human beings. Deep learning based edge detectors have become the mainstream methods for current edge detection.
As the pioneer of contemporary edge detection, HED first introduces the deep supervision mechanism to edge detection and learns multi-scale predictions holistically, which dominates the development of edge detection at that time [liu2019richer, he2022bdcn, huan2021unmixing]. To keep the lightweight of the model, these methods use extremely simple decoders to fuse and compress the multi-scale information into single-channel edges. With the development of deep learning, the encoder and decoder are gradually upgraded iteratively. Puet al. [pu2022edter] replace the encoder with Transformer. Denget al. [deng2018learning] introduce encoder-decoder with symmetric structure into edge detection. But this encoder-decoder structure has remained the same. The encoder extracts multi-scale features including location information and semantic information, and the decoder fuses these features to obtain the final edge map.
In recent years, the development of this structure seems to reach a bottleneck and researchers turn to focus on the optimization of data processing and training strategies [zhou2023uaed, ye2024diffusionedge, li2023kded]. These methods further improve the performance of edge detectors by using complex data processing and training strategies such as data distillation [li2023kded], multi-stage training [pu2022edter], multi-step inference [ye2024diffusionedge], auxiliary model [fu2023practical], etc.
@@ 2.1 The Encoder of Edge Detection
The pioneer of contemporary edge detection is HED [xie2017holistically], which extracts features using VGG [simonyan2014very], the state-of-the-art feature extraction network at the time, and fuses multi-scale features with several convolutions and upsampling to get the final edge. Although rough, HED fuses local and global information for the first time, which makes it perform much better than other detectors based on local information and achieves the highest accuracy. Inspired by HED, a series of excellent edge detectors [liu2019richer, he2022bdcn, huan2021unmixing] have emerged. For fair comparison, there is a tacit agreement to use VGG to extract features and make the decoder as lightweight as possible.
With the development of deep learning, stubborn selection of VGG to extract features is not only not conducive to further improvement of model accuracy, but also difficult to flexibly adjust the size of the model. Therefore, researchers begin to explore new encoders. On the one hand, they explore more powerful encoders to obtain higher accuracy edges. Both RCF [liu2019richer] and BDCN [he2022bdcn] try to directly use ResNet [he2016deep], a more advanced encoder to improve the performance of their models. EDTER [pu2022edter] introduces Transformer into edge detection for the first time, which brings huge computation, but also improves the accuracy by leaps and bounds. The important innovation of DiffusionEdge [ye2024diffusionedge] is to introduce Diffusion to edge detection, but its encoder is also updated with the latest SWin-Transformer. On the other hand, more efficient encoders are designed to obtain lightweight models, FINED [wibisono2021fined] only uses the first few convolutions of VGG to avoid a lot of computation, and LDC [soria2022ldc] utilizes channel separable convolutions and a smaller number of channels to design a more lightweight encoder. PiDiNet [su2021pixel] combines channel separable convolutions with handcrafted features to obtain high-quality edges with a lighter weight model.
Edge detection requires location information and semantic information. In previous works [liu2019richer, pu2022edter, zhou2023uaed], both are extracted by the same encoder. This may seem to simplify the design of the model, but shallow features need to maintain high resolution to ensure the accuracy of location information, while keep a sufficient number of channels to provide rich information cues for deep features. This leads to an inevitable increase in model parameters and computational cost. This problem is exacerbated after Transformer is introduced to edge detection [pu2022edter, ye2024diffusionedge], which results in tens of times more parameters compared to previous CNN-based methods [liu2019richer, he2022bdcn].
@@ 2.2 The Decoder of Edge Detection
To the best of our knowledge, the two most commonly used decoders in edge detection are the HED-style decoder [xie2017holistically, he2022bdcn, xuan2022fcl] and the UNet-style decoder [deng2018learning, deng2020deep]. The HED-style decoder fuses and compresses the multi-scale features into a single-channel edge map using only simple upsampling and $1\times 1$ convolution. The HED-style decoder can obtain the location information in the shallow features and the semantic information in the deep features at the same time. However, shallow features will contain a lot of noise due to the lack of semantic information, and semantic features will introduce location errors due to excessive downsampling. The existence of these two kinds of errors seriously affects the quality of the generated edge map. The UNet-style decoder gradually fuses high-level semantic information with shallow features, and compresses the fused shallowest features into a single-channel edge map. Therefore, the adverse effects of location errors in the semantic information on the edge map can be alleviated. However, the close contact between shallow features and the final result inevitably mixes noise into the generated edge map.
@@ 2.3 Training Strategies for edge detection
UAED [zhou2023uaed] studies the subjectivity and ambiguity of different annotations by estimating the uncertainty of the data. This not only requires the dataset to have multiple sets of labels, but also requires additional processing of these labels. KDED [li2023kded] and PEdge [fu2023practical] eliminate the adverse effects caused by subjectivity and ambiguity of annotations by distillation and self-distillation, which make the training process more complex. PiDiNet [su2021pixel] and PiDiNeXt [li2023pidinext] achieve efficient inference through reparameterization, and EDTER [pu2022edter] achieves global context modeling and local refinement through multi-stage training, respectively. DiffusionEdge [ye2024diffusionedge] introduces Diffusion to the edge detection task and transforms edge detection into a multi-step denoising process.
@@ 3 A NEW BASELINE FOR EDGE DETECTION
[…]
@@ 3.1 Location Feature Encoder
In previous works [liu2019richer, he2022bdcn, xuan2022fcl], multi-scale features are usually extracted by the same encoder. As a result, the shallow module is not only used to generate location features, but also used to provide cues for higher-level modules. For the shallow feature $L_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$, to ensure the accuracy of the location information, the resolution $H_{i}\times W_{i}$ must be large enough. At the same time, the number of channels $C_{i}$ cannot be reduced to keep rich cues for semantic information. So $L_{i}$ is usually a feature with heavy parameters. In addition, the problem of gradient instability forces researchers to employ additional processing such as regularization [ioffe2015batch] and residual connection [he2016deep], even in shallow modules. It makes the computation of shallow features more complex.
We propose an independent location feature encoder, a shallow network. The location feature encoder contains two vanilla Convolution blocks, and each vanilla convolution block consists of a $3\times 3$ convolution module and a ReLU module. To ensure the accuracy of the location information, the feature resolution in the first vanilla Convolution block is consistent with the input image. A down-sampling module is inserted between the two convolution blocks to generate two-scale location features, which can reduce the scale gap between location features and semantic features. Consistent with VGG16 [simonyan2014very], the number of channels doubles when the feature resolution is halved. Since providing cues for semantic information is no longer considered, the features in Location Feature Encoder is compressed to 16 channels, which is much smaller than existing models such as VGG16 [simonyan2014very] (64 channels) and ResNet50 [he2016deep] (64 channels). Compared to VGG16 and ResNet50, our Location Feature Encoder not only has fewer modules, but also has fewer channels per module. Therefore, even though the features keep high resolution, the computational cost remains low.
@@ 3.2 Semantic Feature Encoder
The cascaded residual block is actually an important component of the Transformer and together with the Self-Attention module contributes to the success of the Transformer [liu2021swin]. We intended to use Transformer to capture the semantic features needed for edge detection, but the massive computational cost of Self-Attention (SA) prevents us, even if the image is downsampled before being fed into the semantic feature encoder. Inspired by CAFormer [yu2023metaformer], we adopt a CNN-Transformer hybrid model that uses CNN in the first two stages with large resolution and SA only on the low-resolution features in the last stage. To further reduce the computational cost, the CNN in the first two stages is replaced by channel separable convolution, which is consistent with CAFormer [yu2023metaformer].
@@ 3.4 Network Training
We adopt the annotator-robust loss proposed in RCF [liu2019richer] for all generated edge maps. It can be formulated as
@@
where p denotes the final edge prediction and y is the ground truth. $\alpha=\lvert Y_{+}\rvert/\lvert Y\rvert$ and $\beta=\lambda\cdot\lvert Y_{-}\rvert/\lvert Y\rvert$. $\lvert Y_{+}\rvert$ and $\lvert Y_{-}\rvert$ are used to represents the number of edge and non-edge, separately. $\lambda$ controls the weight of positive over negative samples. $\lvert Y\rvert=\lvert Y_{+}\rvert+\lvert Y_{-}\rvert$. Due to the inconsistency of annotations among different annotators, a threshold $\gamma$ is introduced for loss computation [soria2023dexined_ext]. For pixel i, it will be regarded as edge if the ground truth $y_{i}$ is more than $\gamma$, and we define the ground truth $y_{i}=1$. Pixel i will be regarded as no-edge if $y_{i}=0$. $\lambda$ controls the weight of edge over non-edge.
@@ 4.1 Datasets
We evaluate the proposed NBED on three widely used datasets BSDS500, NYUDv2, and BIPED. BSDS500 [arbelaez2010contour] is by far the most widely used dataset in edge detection, which includes 200 training images, 100 validation images, and 200 test images. Four to nine annotated maps are associated with each image. The final edge ground truth is computed by taking their average. NBED are trained on training and validation images, and tested on the test images. NYUDv2 [silberman2012indoor] is an indoor scene semantic segmentation dataset, the edge ground truth is generated from segmentation maps. it contains 1449 groups of carefully annotated RGB and depth images, and each group image has one annotation result. We use 795 images to train the model and evaluate the model on the rest images. BIPED [poma2020dense, soria2023dexined_ext] contains 250 carefully annotated high-resolution Barcelona Street View images. There are 200 images for training and validation, and 50 images for testing. All images are carefully annotated at single-pixel width by experts in the computer vision field.
The data augmentation strategy during training is shown in Table 2. The data augmentation strategy of BSDS is consistent with LPCB [deng2018learning]. The data augmentation strategy of NYUDv2 follows RCF [liu2019richer], with the only difference that the augmentation is randomly cropped to $400\times 400$ for batch calculation. The data augmentation strategy of BIPED is consistent with DexiNed [poma2020dense].
@@
BSDS F ($4\times$), R ($25\times$) $481\times 321$
NYUDv2 F ($2\times$), S ($3\times$), R ($4\times$) $400\times 400$
BIPED F ($2\times$), S ($3\times$), R ($16\times$), G ($3\times$) $400\times 400$
@@ 4.2 Implementation details
Due to local correlation, the edges obtained by deep learning-based methods are edge probability maps, even after NMS processing. So we need to choose a threshold to binarize the edges. There are two options for the threshold used to binarize the edges. One is to find an optimal threshold value for each image (OIS), and another is to use the same optimal threshold for the whole dataset (ODS). The maximum tolerance allowed for correct matches between edge predictions and ground truth annotations is set to 0.0075 for BSDS500 and BIPED, and 0.011 for NYUDv2. For more experimental details, please refer to previous works [huan2021unmixing, poma2020dense].
@@ 4.3 Comparison with state of the arts
Methods Pub.’Year $S$ MS Param. (M) GFLOPs
@@
Canny [canny1986computational] PAMI’86 0.611 0.676 - - - -
gPb-UCM [arbelaez2010contour] PAMI’10 0.729 0.755 - - - -
SCG [ren2012discriminatively] NeurIPS’12 0.739 0.758 - - - -
SE [dollar2014fast] PAMI’14 0.743 0.764 - - - -
OEF [hallman2015oriented] CVPR’15 0.746 0.770 - - - -
DeepEdge [bertasius2015deepedge] CVPR’15 0.753 0.772 - - - -
DeepContour [shen2015deepcontour] CVPR’15 0.757 0.776 - - - -
Deep Boundary [kokkinos2015pushing] ICLR’15 0.789 0.811 0.803 0.820 - -
CEDN [yang2016object] CVPR’16 0.788 0.804 - - - -
RDS [liu2016learning] CVPR’16 0.792 0.810 - - - -
AMH-Net [xu2017learning] NeurIPS’17 0.798 0.829 - - - -
HED [xie2017holistically] PAMI’17 0.788 0.808 - - 14.6 57.5
CED [wang2017deep] CVPR’17 0.803 0.820 - - 21.8 -
LPCB [deng2018learning] ECCV’18 0.800 0.816 - - - -
RCF [liu2019richer] PAMI’19 0.798 0.815 - - 14.8 75.3
DSDC [deng2020deep] ACMMM’20 0.802 0.817 - - - -
LDC [deng2021learning] ACMMM’21 0.799 0.816 - - - -
PiDiNet [su2021pixel] ICCV’21 0.789 0.803 - - 0.71 11.6
BDCN [he2022bdcn] PAMI’22 0.806 0.826 - - 16.3 103.4
FCL [xuan2022fcl] NN’22 0.807 0.822 0.816 0.833 16.5 134.4
EDTER [pu2022edter] CVPR’22 0.824 0.841 0.840 0.858 468.8 802.3
UAED [zhou2023uaed] CVPR’23 0.829 0.847 0.837 0.855 72.5 72.8
DiffusionEdge [ye2024diffusionedge] AAAI’24 0.834 0.848 - - 225 494
@@
We compare our model with traditional detectors (Canny [canny1986computational], SCG [ren2012discriminatively], OEF [hallman2015oriented], etc), CNN-based detectors (RCF [liu2019richer], BDCN [he2022bdcn], UAED [zhou2023uaed], etc) and the transformer-based detectors (EDTER [pu2022edter] and DiffusionEdge [ye2024diffusionedge]) on BSDS500. Quantitative results are shown in Table 3. By training on the trainval set of BSDS500, our NBED achieves the F-measure ODS of 0.834 with single-scale testing and obtains 0.840 with multi-scale testing, which is superior to all the state-of-the-art edge detectors.
It’s worth mentioning that NBED is a vanilla encoder-decoder with neither a complex training strategy nor a huge model size, which stands out among recent high performance edge detectors. Among the detectors with similar performance to our NBED, EDTER [pu2022edter] not only has a model size of 11 times larger than NBED, but also requires multi-stage training. Nonetheless, the single-scale ODS of EDTER is still 1.4% lower than that of NBED. The number of parameters of UAED [zhou2023uaed] is about twice that of NBED, and it requires multi-label data. The single-scale ODS of UAED in this case is 0.9% lower than that of NBED. DiffusionEdge [ye2024diffusionedge] appears to perform competitively, with a single-scale ODS 0.4% lower than NBED’s, and DiffusionEdge not only relies on large-scale models, but also requires multi-step sampling during inference, which is very inefficient. In addition to these methods with complex training strategies, the top performing vanilla encoder-decoder is FCL [xuan2022fcl] with a single-scale ODS of 0.807, which is more than 3% lower than NBED and is no longer a contender. Some qualitative results are shown in Fig. 2. We observe that the proposed NBED shows a clear advantage in prediction quality, both crisp and accurate.